# What makes math problems hard for reinforcement learning: a case study

### v1

@@ 4.3. Application to the Andrews–Curtis Conjecture
We also need to choose a horizon length $T$ over which the cumulative return is calculated (c.f. Subsection 4.1). In reinforcement learning, training on tasks with long horizons is usually harder than on those with short horizons. This is because in long horizon tasks it is often more difficult to figure out which actions are responsible for the eventual outcomes, a problem known as the credit assignment problem. There is also more uncertainty in the returns (high variance), which can slow down convergence and destabilize the learning process. The size of the exploration space also becomes a critical issue. With longer horizons, the agent needs to explore a more diverse set of actions to learn about long-term consequences without getting stuck in suboptimal regions. However, it does not mean that we can not train well-performing agents for long-horizon tasks; rather, it indicates that with longer horizons we may need significantly stronger computational power and extended training periods. [^8] Another option for improving performance in long-horizon tasks is to provide good intermediate rewards; unfortunately, this is rather hard in the present context of the AC conjecture.
The horizon length $T$ is an important hyperparameter in the context of our problem as any presentation that requires a sequence of AC-moves of length greater than the horizon length will necessarily remain unsolved by PPO. On the other hand, choosing a horizon length that is too long can significantly slow down the training process. Due to the limited amount of hardware available, we mostly experimented with considerably smaller values for $T$, i.e. $T=200$ and $T=400$. With these values, PPO could solve, respectively, $431$ and $402$ presentations out of the $1190$ Miller–Schupp presentations of the initial state distribution. [^9] We also explored the value $T=2000$. However, we found it much slower to train due to the reasons described above. We could only solve $219$ presentations of the initial state distribution in this case. This training run had not converged, and we expect that with more computational power and extended training periods, it will be worthwhile to experiment with larger values of $T$, perhaps helping us solve even more presentations than greedy search. In each case, these presentations formed a subset of the presentations solved by the greedy search.
@@
In 4(b), we give a scatterplot of the path lengths discovered by PPO and greedy search for all of the presentations solved by PPO. We note that in many cases, PPO found shorter paths compared to the greedy search. This is expected as PPO learns to improve its strategy during the training, discovering shorter paths for presentations it may have already solved. The scatterplot shows the shortest paths discovered by PPO for each presentation. We also note that in many cases, PPO found longer paths than greedy search. This shows that our specific run exhibits a suboptimal policy. It could perhaps be improved by performing more hyperparameter tuning on the training process.
@@ 5. The Cure: New Algorithms
In previous sections we explained from a variety of different perspectives that the Andrews–Curtis conjecture is a good example of a mathematical problem where the length of a solution can be much greater than the length of the initial presentation, in some cases with purely analytical lower bounds that are hyperexponential in the size of the input. In particular, we saw that small increases in presentation length under 20 quickly lead to solution lengths in the range of hundreds and higher, quickly exceeding the number of moves in the longest game of chess.
If solving a mathematical problem required finding a path of length $L$, say with $L=10^{6}$, an RL agent would be pretty much out of luck under circumstances of a typical hard search problem, where the number of successful paths is exponentially suppressed by $L$. The good news is that in mathematics —and in many other domains— such hard search problems never come in isolation. Rather, there is a distribution of problems such that generic instances are “easy” and a small fraction is “hard.” Learning this distribution for smaller values of $L$ contains the crucial information for solving new cases at the next increment of $L$.
@@ 5.1. Supermoves
In automated reasoning or search problems where the minimal length solution has a theoretical lower bound that by far exceeds computational capabilities, it is clear that direct approach with fixed size steps is not going to succeed, unless the problem is easy and a large fraction of long paths meets the desired criteria. In order to reach extraordinary path lengths, one must allow progressively longer sequences of elementary moves to be added to the action space. Although this general strategy seems unavoidable in problems like the AC conjecture, it leads to many practical questions. For example, what should be the selection criteria for such “supermoves”? And, how often should they be added to the action space?
In the context of the AC conjecture, a good example of such supermoves are the “elementary M-transformations” [BurnsI, BurnsII]. These transformations trivialize $\text{AK}(2)$ in just two steps, even though this presentation is known to admit the shortest AC trivialization path of length 14. A downside of elementary M-transformations, though, is that they are infinite in number, which complicates their application in classical search techniques.
In our study, we explored the idea of identifying AC supermoves by selecting some frequently occurring subsequences of AC-moves in the paths discovered by Proximal Policy Optimization (PPO). By extending the action space $A$ of the Markov Decision Process (MDP) with these subsequences and checking whether this enhanced action space helps our agent discover shorter paths of trivialization, we learned a few useful lessons:
@@
5.2. The anatomy of success
While supermoves clearly need to be a part of the solution in hard problems like the AC conjecture, much of the success depends on the criteria for selecting them. Here, we advocate for a dynamic approach where the network itself learns the criteria for selecting supermoves, in addition to the best ways to implement them. One realization of this approach could be a multi-agent model, where one network is learning to play the game and the other is learning the rules for changing the action space (adding and removing supermoves). We hope that future iterations of this strategy can lead to AI systems that can ‘learn how to learn’ dynamically by making both algorithmic and architectural changes through collecting the information about hard instances. [^12] Here, by self-improving AI systems we mean algorithms that have the ability to “interpolate” between off-the-shelf algorithms such as A2C and TRPO, as well as a myriad of custom algorithms that do not even have a name. Clearly, this level of technology is not presently available, and one of the key points of this section is that developing such systems should be based on the hardest instances the agent encounters.
@@ 5.2. The anatomy of success
Specifically, suppose $N$ is one of the characteristics of either the algorithm or the architecture that has non-trivial impact on performance. In practice, there can be several such parameters, but for simplicity we explain the idea as if there is only one. [^13] Analyzing a multi-dimensional landscape is generally a good idea, as it can yield better performance improvements, but the price to pay is that changes of different characteristics become correlated, as illustrated in Figure 6. Then, from the practical standpoint, a natural notion of hardness is such that hard instances are defined to be those which the model can solve at the current setting of $N$ and not with the lower value of the resource $N$. In addition, in search problems we include the length of the path in the notion of hardness, i.e. select a subset of the instances that the model could solve through especially long paths. Note, by the very nature of the search problem we are interested in, there can not be too many such hard instances at each step of increasing $N$, for otherwise the problem would be easy, not hard. Collecting the information about the hardest instances at each increment in $N$ can be used to select supermoves, e.g. as subsequences of the sequences of moves that solve the hard instances. Algorithm 3 provides one particular realization of this idea.
@@
In the context of the AC conjecture, examples of the metric $N$ can be the horizon length or the number of interactions with the environment. As Figure 5 illustrates, increasing the number of environment interactions leads to a larger number of non-trivial presentations from the Miller–Schupp series being solved (i.e. AC-trivialized) by our RL agent. Moreover, the length of the AC trivialization path also grows for some of the solutions (but not all). Therefore, in order to implement the program outlined above in the specific context of the AC conjecture, we can focus on the longest AC trivialization paths that the model is able to find at each value of $N$.
@@
Collecting this information in the process of training can be used to introduce (and remove) supermoves to the action space in a dynamic fashion. There can be many different implementations of this idea that we plan to explore more fully elsewhere. For example, one can consider selecting all or subset of the longest AC trivialization paths that the model finds at each $N$. Out of those, in each case, one can consider selecting the entire trivialization path or a subset, randomly or non-randomly. Alternatively, one can compare the longest trivialization paths at several (consecutive) values of $N$ and choose subsequences of moves that are shared by several long trivialization paths at different $N$.
For example, if $N$ denotes the number of interactions with the environment, we did a few preliminary experiments with the dataset of Section 4 and several different seed values. To illustrate the effects of stochasticity, let us consider $N=8\times 10^{7}$. The agents with five different seed values were able to solve 354, 337, 330, 328, and 323 presentations, respectively. And their average, $334.4$, is shown in Figure 5. Many of these AC presentations can be solved at the earlier stage, with $N=4\times 10^{7}$ or less. If in the definition of hard instances we require that they are solved by all five agents, there are only 5 presentations total. On the other hand, if we require that they are solved by any of the 5 agents, the number goes up to 36.
@@
Moreover, not surprisingly, the 5 presentations solved by all 5 agents have considerably shorter path lengths (47, 31, 22, 14, and 13) compared to path lengths of the 36 presentations illustrated on the left panel of Figure 7 that go up to $200$. Both 5 presentations solved by all agents and 36 presentations solved by at least one of the agents provide viable options for defining hard instances and, in turn, selecting supermoves. However, they lead to qualitatively different results. For example, all 5 presentations solved by all 5 agents are solved at a smaller value of $N$ when required to be solved by only one of the agents. More importantly, they have very different anatomy, illustrated in Figure 8 and in Figure 9, respectively. By examining the longest trivialization paths of the 36 presentations solved by at least one agent at $N=8\times 10^{7}$, we often see long strings of moves #$\#5$ and #$\#11$, interlaced with moves #$\#3$, #$\#7$, and #$\#9$. These are our top candidates for the supermoves to be added at $N=8\times 10^{7}$. Note that moves #$\#4$ and #$\#8$ are least common in the examples presented in both Figure 8 and Figure 9.
@@
6. Isolated components and neighborhood sizes
6.1. Isolated components
In Subsection 3.2, we explored a greedy approach to finding AC-trivializations of a presentation $\pi$. Specifically, the goal was to construct a sequence of presentations $(\pi_{0},\dots,\pi)$, where $\pi_{0}$ is the trivial presentation, such that each consecutive pair in the sequence is related by an AC-move. Furthermore, at each step $k$, the presentation $\pi_{k}$ was chosen to have the shortest possible length among all presentations connected to $\pi_{k+1}$ via an AC-move. In general, the length of a presentation in an AC-trivialization tends to exceed the length of the original presentation being trivialized. The minimum increase in length across all possible AC-trivializations serves as an invariant of the presentation. We will explore this invariant using concepts from persistent homology.
6.1.1. Formalization
A based graph is a pair $(\Gamma,v_{0})$ consisting of a graph $\Gamma$ and a preferred vertex $v_{0}$ in it. We will often drop $v_{0}$ from the notation. A based subgraph $\Gamma_{n}$ of $\Gamma$, written $\Gamma_{n}\leq\Gamma$, is a subgraph $\Gamma_{n}$ of $\Gamma$ with the same preferred vertex. We say that $\Gamma_{n}$ is full in $\Gamma$ if for any two vertices in $\Gamma_{n}$ joined by an edge in $\Gamma$, the edge is also in $\Gamma_{n}$. A filtration of a based graph $\Gamma$ is a collection
$\Gamma_{0}\leq\Gamma_{1}\leq\Gamma_{2}\leq\dotsb$
of based subgraphs of $\Gamma$ for which each vertex and edge of $\Gamma$ is in $\Gamma_{n}$ for some $n$. We refer to $\Gamma$ as a filtered based graph. If each $\Gamma_{n}$ is full in $\Gamma$ we refer to the filtration as full and notice that full filtrations are equivalent to $\mathbb{N}$ valued functions from the set of vertices of $\Gamma$ sending $v_{0}$ to 00.
Let $\Gamma^{\text{AC(k)}}$ be the graph whose vertices are $k$-balanced presentations, based at the trivial presentation, having an edge between two vertices if there is an AC-move between them. Additionally, $\Gamma^{\text{AC(k)}}$ is equipped with a full filtration obtained from the function sending a vertex to the length of its presentation minus $k$.
Given a filtered based graph $(\Gamma,v_{0})$. The filtration value $\operatorname{Filt}(v)$ of a vertex $v$ is the smallest $n\in\mathbb{N}$ such that $v$ is a vertex in $\Gamma_{n}$. Similarly, its connectivity value $\operatorname{Conn}(v)$ is the smallest $n\in\mathbb{N}$ such that $v$ and $v_{0}$ can be joined by a path in $\Gamma_{n}$ or is set to $\infty$ if such path does not exist in $\Gamma$.
The isolation value of a vertex $v$ in a filtered based graph is defined as
$\operatorname{Isol}(v)=\operatorname{Conn}(v)-\operatorname{Filt}(v),$
a number in ${}\mathbb{N}\cup\{\infty\}$. A vertex is said to be isolated if its isolation value is positive.
We introduce an equivalence relation on isolated vertices. Two belong to the same isolated component if they have the same connectivity value, say $n$, and they can be joined by a path in $\Gamma_{n-1}$. The isolation value of a component is the maximum of the isolation values of its elements.
We can interpret these invariants using the framework of topological data analysis, see for example [carlsson2022tda]. Specifically, the set of isolated components of a based filtered graph $\Gamma$ corresponds to the multiset of bars in the barcode of its reduced persistent 00-homology. Additionally, the isolation value of an isolated component corresponds to the length of its associated bar.
6.1.2. Experimental results
Let $\Gamma^{\ell}$ be the full subgraph of $\Gamma^{\text{AC(2)}}$, with the induced filtration, consisting of all presentations with a connectivity value less than or equal to $\ell$. Explicitly, $\Gamma^{\ell}$ includes all vertices that can be connected to the trivial vertex via paths containing only presentations of length at most $\ell$.
We will denote by $v(\ell)$ and $e(\ell)$ the number of vertices and edges of $\Gamma^{\ell}$. Let us denote by $ic(\ell)_{k}$ the number of isolated components with isolation value $k$. Figure 11 summarize our results for the classic AC-moves whereas Figure 12 does so for their prime version. [^14] For this task we used giotto-TDA version 5.1 [tauzin2021giotto]. Specifically, its binding of the SimplexTree data structure introduced in [boissonnat2014simplex] and implemented in GUDHI [maria2014gudhi].
$\ell$ $v(\ell)$ $e(\ell)$ $ic(\ell,1)$ $ic(\ell,2)$ $ic(\ell,3)$
3 36 72 0 0 0
4 100 248 0 0 0
5 388 1072 0 0 0
6 884 2376 0 0 0
7 3892 10775 0 0 0
8 9172 25675 0 0 0
9 37428 106513 0 0 0
10 84996 239733 0 0 0
11 350356 1002439 4 0 0
12 791140 2251375 16 0 0
13 3238052 9321629 72 4 0
14 7199908 20573343 144 4 0
15 29243812 84391763 508 52 8
16 64623652 185162236 1034 88 20
Figure 11. Classical AC-moves
$\ell$ $v(\ell)$ $e(\ell)$ $ic(\ell,1)$ $ic(\ell,2)$ $ic(\ell,3)$
3 36 40 3 0 0
4 100 152 3 0 0
5 388 712 3 0 0
6 884 1528 3 0 0
7 3892 6984 3 0 0
8 9172 16728 3 0 0
9 37428 69752 3 0 0
10 84996 155752 3 0 0
11 350356 655928 19 0 0
12 791140 1467080 67 0 0
13 3238052 6107112 243 16 0
14 7199908 13414744 483 16 0
15 29243812 55306744 1819 136 32
16 64623652 120824232 3923 208 80
Figure 12. Prime AC-moves
6.2. Neighborhoods
Let us return to our data set of 1190 presentations in the Miller–Schupp series for $n\leq 7$ and $\operatorname{length}(w)\leq 7$. Using the methods described in Subsection 4.2, we trained a PPO agent that successfully solved 417 of these presentations. We will refer to the set of these 417 presentations as PPO-solved and the remaining 773 presentations as PPO-unsolved. Our goal is to analyze the relationship between these labels and the sizes of their respective AC neighborhoods. A presentation is considered to be in the $k$-step neighborhood of another if they can be connected by applying at most $k$ AC-moves.
6.2.1. Experimental results
There are 131 distinct neighborhood sizes in our data. Their basic statistics are
Min Max Mean Median
72,964 89,872 89,532 89,859
A more detailed description of the frequency of values is presented in Figure 13.
The largest neighborhood size accounts for nearly a third of all considered presentations. However, it represents only 2.4% of PPO-solved presentations, while constituting almost half (48.3%) of the PPO-unsolved presentations. For more details, please refer to Figure 14.
As we observed, having a maximal neighborhood size provides significant insight into whether a presentation is labeled as PPO-solved or PPO-unsolved. Additionally, the minimum neighborhood size among PPO-unsolved presentations—89,573—is also quite telling, as 54% of PPO-solved presentations have neighborhood sizes smaller than this value. This percentage can be further improved by considering that the neighborhood sizes of PPO-unsolved presentations are concentrated within three specific bands. Please refer to Figure 15 for more details. We find that 64.3% of PPO-solved presentations fall outside the three bands $[89,575,89,575]$, $[89,715,89,831]$, and $[89,844,89,872]$, which together contain over 99% of PPO-unsolved presentations. By narrowing the last band to $[89,859,89,872]$, these three bands now encompass the neighborhood sizes of over 90% of PPO-unsolved presentations, while their complement includes 77.2% of PPO-solved presentations.
One might expect that enhancing the discriminatory power of $n$-neighborhoods could be achieved by incorporating features beyond their size. We explored two additional types of features, but surprisingly, they only marginally improved the accuracy of PPO-solved/unsolved predictions. The first type was based on node centrality, while the second focused on spectral features of the neighborhood graphs. The latter was particularly intriguing, given the emphasis on Markov processes and the well-known relationship between random walks on graphs and the graph Laplacian.

### v2

@@ 4.3. Application to the MS series
We also need to choose a horizon length $T$ over which the cumulative return is calculated (c.f. Section 4.1). In reinforcement learning, training on tasks with long horizons is usually harder than on those with short horizons. This is because in long horizon tasks it is often more difficult to figure out which actions are responsible for the eventual outcomes, a problem known as the credit assignment problem. There is also more uncertainty in the returns (high variance), which can slow down convergence and destabilize the learning process. The size of the exploration space also becomes a critical issue. With longer horizons, the agent needs to explore a more diverse set of actions to learn about long-term consequences without getting stuck in suboptimal regions. However, it does not mean that we can not train well-performing agents for long-horizon tasks; rather, it indicates that with longer horizons we may need significantly stronger computational power and extended training periods. [^12] Another option for improving performance in long-horizon tasks is to provide good intermediate rewards; unfortunately, this is rather hard in the present context of the AC conjecture.
The horizon length $T$ is an important hyperparameter in the context of our problem as any presentation that requires a sequence of AC-moves of length greater than the horizon length will necessarily remain unsolved by PPO. On the other hand, choosing a horizon length that is too long can significantly slow down the training process. We experimented with two different schedules for horizon length during training:
  1. (i)
    constant horizon length throughout training, and
  2. (ii)
    varying horizon length during training.
The trained agents behave differently in the two cases as we will now describe.
4.4. PPO with constant horizon length: experimental results
When keeping the horizon length fixed during training, we experimented with considerably smaller values, i.e. $T=200$ and $T=400$. Out of the $1190$ Miller–Schupp presentations of the initial state distribution, the best performing PPO agents could solve $431$ and $402$ presentations respectively.  [^13] We also explored the value $T=2000$. However, we found it much slower to train due to the reasons described above. We could only solve $219$ presentations of the initial state distribution in this case. This training run had not converged, and we expect that with more computational power and extended training periods, it will be worthwhile to experiment with larger values of $T$, perhaps helping us solve even more presentations than greedy search. In each case, these presentations formed a subset of the presentations solved by the greedy search (Figure 5).
(a) Distribution versus $n$
(b) Distribution versus length
Figure 5. A comparison of three algorithms —breadth-first search, greedy search, and Proximal Policy Optimization (PPO) with small and constant horizon length during training — that we used to search through the space of balanced presentations. The number of presentations of the Miller–Schupp series, $\text{MS}(n,w)$, solved by an algorithm is given on the vertical axis. We compare the performance as a function of $n$ (above) and the length of the presentation (below). Note that varying horizon length during training (not depicted in this figure) helped PPO solve presentations that greedy search could not solve.
@@ 4.4. PPO with constant horizon length: experimental results
In Figure 6(b), we give a scatterplot of the path lengths discovered by PPO and greedy search for all of the presentations solved by PPO. We note that in many cases, PPO found shorter paths compared to the greedy search. This is expected as PPO learns to improve its strategy during the training, discovering shorter paths for presentations it may have already solved. The scatterplot shows the shortest paths discovered by PPO for each presentation. We also note that in many cases, PPO found longer paths than greedy search. This shows that our specific run exhibits a suboptimal policy. It could perhaps be improved by performing more hyperparameter tuning on the training process.
@@
4.5. New Miller–Schupp trivializations through variable horizon length
Keeping the horizon length fixed either fails to solve hard presentations if the horizon length is too small, or requires significant computational resources for training to converge if the horizon length is too large. Therefore, we experimented with varying horizon length during training.
We trained a PPO agent for 100 million environment interactions, with the same hyperparameters as in the previous section; only varying the horizon length over the course of training. In the beginning, the horizon length was set to $200$; it was increased to $400$ after 10 million environment interactions, then to $800$ after 25 million environment interactions, and finally to $1200$ after 50 million environment interactions. This schedule allowed the PPO agent to first solve easier presentations, avoiding the challenges of long-horizon tasks. As the agent’s capabilities improved, the gradual increase in horizon length enabled the agent to tackle more complex Miller-Schupp presentations.
The best-performing PPO agent solved two new presentations of the Miller-Schupp series that the greedy search had failed to solve.
Theorem 10.
The following potential counterexamples belonging to the Miller–Schupp series are AC-trivial:
$\langle x,y\mid x^{-1}y^{3}xy^{-4},\ x^{-1}y^{-1}xy^{-3}x^{-1}y^{-1}\rangle,$
$\langle x,y\mid x^{-1}y^{4}xy^{-5},\ x^{-1}yxy^{-1}x^{-1}y^{-3}\rangle.$
For these presentations, the AC paths discovered by the agent had lengths of $205$ and $1085$, respectively. These paths contained subsequences of moves that acted trivially. We removed these subsequences, obtaining paths of lengths $195$ and $381$ respectively. The initial lengths of the two presentations are $17$ and $19$, which increased to the maximum values of $29$ and $30$ before the presentations were trivialized.
With the following definitions of AC moves,
the length 195 path for the first of the two presentations above is as follows.
$h_{2}\cdot h_{8}\cdot h_{10}\cdot h_{4}\cdot h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{4}$
$h_{0}\cdot h_{6}\cdot h_{2}\cdot h_{8}\cdot h_{8}\cdot h_{10}\cdot h_{4}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{4}$
$h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{4}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}$
$h_{4}\cdot h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{4}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}$
$h_{6}\cdot h_{4}\cdot h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{4}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{2}\cdot h_{8}$
$h_{6}\cdot h_{6}\cdot h_{4}\cdot h_{10}\cdot h_{10}\cdot h_{10}\cdot h_{1}\cdot h_{7}\cdot h_{5}\cdot h_{5}\cdot h_{5}\cdot h_{11}\cdot h_{9}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{4}\cdot h_{2}$
$h_{8}\cdot h_{6}\cdot h_{4}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{4}\cdot h_{10}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{2}$
$h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{8}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{4}\cdot h_{10}\cdot h_{8}\cdot h_{6}\cdot h_{2}\cdot h_{8}\cdot h_{6}$
$h_{6}\cdot h_{8}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{2}\cdot h_{8}\cdot h_{6}\cdot h_{6}\cdot h_{8}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{6}$
$h_{8}\cdot h_{8}\cdot h_{6}\cdot h_{8}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{6}\cdot h_{8}\cdot h_{8}\cdot h_{0}\cdot h_{6}\cdot h_{8}\cdot h_{6}\cdot h_{8}\cdot h_{10}\cdot h_{0}$
$h_{6}\cdot h_{8}\cdot h_{6}\cdot h_{8}\cdot h_{8}\cdot h_{1}\cdot h_{7}\cdot h_{5}\cdot h_{11}\cdot h_{9}\cdot h_{3}\cdot h_{11}\cdot h_{9}\cdot h_{11}\cdot h_{8}\cdot h_{0}\cdot h_{6}\cdot h_{3}\cdot h_{11}.$
This sequence should be read from left to right; first apply $h_{2}$, then $h_{8}$, and so forth. The length 381 path is presented in Appendix B.  [^16] An interested reader may verify the correctness of these paths through a Jupyter Notebook available at the following link for the associated GitHub repository, https://github.com/shehper/AC-Solver.
@@ 5. The cure: new algorithms
In previous sections we explained from a variety of different perspectives that the Andrews–Curtis conjecture is a good example of a mathematical problem where the length of a solution can be much greater than the length of the initial presentation, in some cases with purely analytical lower bounds that are superexponential in the size of the input. In particular, we saw that small increases in presentation length under 20 quickly lead to solution lengths in the range of hundreds and higher, quickly exceeding the number of moves in the longest game of chess.
If solving a mathematical problem required finding a path of length $L$, say with $L=10^{6}$, an RL agent would be pretty much out of luck under circumstances of a typical hard search problem, where the number of successful paths is exponentially suppressed by $L$. The good news is that in mathematics — and in many other domains — such hard search problems never come in isolation. Rather, there is a distribution of problems such that generic instances are “easy” and a small fraction is “hard.” Learning this distribution for smaller values of $L$ contains crucial information that one can use to solve new cases with the larger value of $L$.
@@ 5.1. Supermoves
In automated reasoning or search problems where the minimal length solution has a theoretical lower bound that by far exceeds computational capabilities, it is clear that direct approach with fixed size steps is not going to succeed, unless the problem is easy and a large fraction of long paths meets the desired criteria. In order to reach extraordinary path lengths, one must allow progressively longer sequences of elementary moves to be added to the action space. [^17] In the AI approaches to theorem proving, this can be compared to constructing proofs of Lemmas and subgoals, which then can be used as building blocks in larger proofs [yang2019coqgym, bansal2019holist, polu2020generative, polu2022formal]. Although this general strategy seems unavoidable in problems like the AC conjecture, it leads to many practical questions. For example, what should be the selection criteria for such “supermoves”? And, how often should they be added to the action space?
In the context of the AC conjecture, a good example of such supermoves are the “elementary M-transformations” [BurnsI, BurnsII]. These transformations trivialize $\text{AK}(2)$ in just two steps, even though this presentation is known to admit the shortest AC trivialization path of length 14. A downside of elementary M-transformations, though, is that they are infinite in number, which complicates their application in classical search techniques; rather, we need a finite action space with a relatively small number of supermoves.
In our study, we explored the simplest version of identifying AC supermoves by selecting some frequently occurring subsequences of AC-moves in the paths discovered by Proximal Policy Optimization (PPO). By extending the action space $A$ of the Markov Decision Process (MDP) with these subsequences and checking whether this enhanced action space helps our agent discover shorter paths of trivialization, we learned a few useful lessons:
@@
Figure 8. A schematic representation of a model with an adaptive action space.
To summarize, the dynamic way of introducing supermoves is crucial for hard search problems like the AC conjecture, and the algorithm (or, policy) for making such adaptive changes to the action space is, in turn, crucial to the performance. Schematically, such adaptive changes to the action space can be illustrated by a diagram in Figure 8, which replaces the traditional RL cycle illustrated in Figure 4. In the rest of this section we discuss several ideas on the implementation of adaptive action spaces motivated by the environment of the AC conjecture. They all are based on the notion of “hardness” which plays an important role in the following.
Figure 9. Depending on the problem, the distribution of new hard instances that are found at each iteration of the model upgrades can vary. This distribution, in turn, determines the update schedule of the action space. The action space of the upgrade $N+1$ is determined by the new hard instances found at the stage $N$ and earlier.
5.2. Adaptive action space and the anatomy of success
While supermoves clearly need to be a part of the solution in hard problems like the AC conjecture, much of the success depends on the criteria for selecting them. Here, we advocate for an adaptive approach where the model itself learns the criteria for selecting supermoves, in addition to the best ways to implement them. One realization of this approach could be a multi-agent model, where one network is learning to play the game while the other is learning the rules for changing the action space (adding and removing supermoves), cf. Figure 8. We hope that future iterations of this strategy can lead to AI systems that can ‘learn how to learn’ dynamically by making both algorithmic and architectural changes through collecting the information about hard instances. [^18] Here, by self-improving AI systems we mean algorithms that have the ability to “interpolate” between off-the-shelf algorithms such as A2C and TRPO, as well as a myriad of custom algorithms that do not even have a name. Clearly, this level of technology is not presently available (see, however, [lu2022discovered] for an interesting work in this direction), and one of the key points of this section is that developing such systems should be based on the hardest instances the agent encounters. In other words, we propose AI system to select supermoves based on the notion of “hardness” that also can be learned.
@@ 5.2. Adaptive action space and the anatomy of success
For concreteness, suppose $N$ is one of the characteristics of either the algorithm or the architecture that has non-trivial impact on performance. In practice, there can be several such parameters, but for simplicity we explain the idea as if there is only one. [^19] Analyzing a multi-dimensional landscape is generally a good idea, as it can yield better performance improvements, but the price to pay is that changes of different characteristics become correlated, as illustrated in Figure 10. Thus, in studying scaling laws one typically looks at performance as a function of $a)$ the number of model parameters (depth and width), $b)$ training time, and $c)$ the size of the data. We can pick $N$ to be any of these, and in the context of the AC conjecture the number of environment interactions (training time) is a natural candidate that we consider below when we need a concrete example. As Figure 10 illustrates, if $N$ represents incremental upgrades in the number of environment interactions, the model of Section 4 is able to expand to larger horizon lengths.
Now, we can define hard instances as those which the model is able solve at the current setting of $N$ and not with the lower value of the resource $N$. In addition, in search problems we include the length of the path in the notion of hardness, i.e. select a subset of the instances that the model could solve through especially long paths. Note, by the very nature of the search problem we are interested in, there can not be too many such hard instances at each step of increasing $N$, for otherwise the problem would be easy, not hard. Collecting the information about the hardest instances at each increment in $N$ can be used to select supermoves, e.g. one can choose the latter to be subsequences of the sequences of moves that solve the hard instances. Algorithm 3 provides one particular realization of this idea.
@@
The reason to include the length of the path in the definition of hardness is well illustrated by the AC conjecture. In this case, increasing the number of environment interactions, $N$, leads to a larger number of non-trivial presentations from the Miller–Schupp series being solved (i.e. AC-trivialized) by our RL agent, see Figure 7. Moreover, the length of the AC trivialization path also grows for some of the solutions (but not all). Therefore, in order to implement the program outlined above in the specific context of the AC conjecture, we can focus on the longest AC trivialization paths that the model is able to find at each value of $N$.
@@
Collecting this information in the process of training can be used to introduce (and remove) supermoves to the action space in an adaptive fashion. There can be many different implementations of this idea that we plan to explore more fully elsewhere. For example, one can consider selecting all or subset of the longest AC trivialization paths that the model finds at each $N$. Out of those, in each case, one can consider selecting the entire trivialization path or a subset, randomly or non-randomly. Alternatively, one can compare the longest trivialization paths at several (consecutive) values of $N$ and choose subsequences of moves that are shared by several long trivialization paths at different $N$.
Below we briefly examine the structure of the longest (i.e. hard) AC trivializations found by our RL agent with several different seed values. First, let us illustrate the effects of stochasticity and consider $N=8\times 10^{7}$. The agents with five different seed values were able to solve 354, 337, 330, 328, and 323 presentations, respectively. And their average, $334.4$, is shown in Figure 7. Many of these AC presentations can be solved at the earlier stage, with $N=4\times 10^{7}$ or less. If in the definition of hard instances we require that they are solved by all five agents, there are only 5 presentations total. On the other hand, if we require that they are solved by any of the 5 agents, the number goes up to 36.
@@
Moreover, not surprisingly, the 5 presentations solved by all 5 agents have considerably shorter path lengths (47, 31, 22, 14, and 13) compared to path lengths of the 36 presentations illustrated on the left panel of Figure 11 that go up to $200$. Both 5 presentations solved by all agents and 36 presentations solved by at least one of the agents provide viable options for defining hard instances and, in turn, selecting supermoves. However, they lead to qualitatively different results. For example, all 5 presentations solved by all 5 agents are solved at a smaller value of $N$ when required to be solved by only one of the agents. More importantly, they have very different anatomy, illustrated in Figure 12 and in Figure 13, respectively. By examining the longest trivialization paths of the 36 presentations solved by at least one agent at $N=8\times 10^{7}$, we often see long strings of moves #$\#5$ and #$\#11$, interlaced with moves #$\#3$, #$\#7$, and #$\#9$. These are our top candidates for the supermoves to be added at $N=8\times 10^{7}$, if $N$ was allowed to gradually increase in an adaptive system, as in Figure 8. Note that moves #$\#4$ and #$\#8$ are least common in the examples presented in both Figure 12 and Figure 13.